OpenStudy (anonymous):

Which of the following is an equivalent form of the compound inequality −22 > −5x − 7 −3? −5x − 7 < −22 and −5x − 7 −3 −5x − 7 > −22 and −5x − 7 −3 −5x > −22 and −7 −3 −5x − 7 < −22 and −5x − 7 −3

OpenStudy (anonymous):

Is the problem written correctly ? Is there supposed to be another inequality sign.....in between the -7 and the -3 ?

OpenStudy (anonymous):

yea its greater than or equal too

OpenStudy (anonymous):

> or equal too

OpenStudy (anonymous):

-22 > -5x - 7 -5x - 7 < -22 (when you switch sides, the inequality sign changes) -5x - 7y >= -3 It is either gonna be a or d. I don't know for sure because the answer choices for the second part of the inequality is not clear. If you could write the complete answer choices for a and d, then I can tell you for sure.

OpenStudy (anonymous):

for A through C its > or equal too and D is < or equal too

OpenStudy (anonymous):

The answer is A :)

OpenStudy (anonymous):

Thankyou :D

OpenStudy (anonymous):

@kelliegirl33 can you help me with some other ones

OpenStudy (anonymous):

let me see what you got

OpenStudy (anonymous):

Which of the following is the correct graph of the solution to the inequality −8 −5x + 2 > −38? Question 2 (Multiple Choice Worth 2 points) Which of the following is the correct graph of the compound inequality 4p + 1 > −11 or 6p + 3 < 39? Question 3 (Multiple Choice Worth 2 points) Which of the following is the correct graph of the compound inequality n > 5 and n < 3?

OpenStudy (anonymous):

number one right after is < or equal too

OpenStudy (anonymous):

are these graphed on a number line ?

OpenStudy (anonymous):

Do you need a number line graph ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

for all three

OpenStudy (anonymous):

ok...let me get to work :)

OpenStudy (anonymous):

1.)-8 <= -5x + 2 -8 - 2 <= -5x -10 <= -5x -10/-5 => x 2 => x -5x + 2 > - 38 -5x > -38 - 2 -5x > -40 x < -40/-5 x < 8 graph : closed circle on 2 and open circle on 8. shading in between the two numbers. 2.) 4p + 1 < -11 4p < -11 - 1 4p < -12 p < - 3 6p + 3 < 39 6p < 39 - 3 6p < 36 p < 6 graph : I am not sure about this one...open circle on -3 and open circle on 6. Start from the 6 and shade to the left past -3. It is double shading when you get to the -3. n > 5 n < 3 graph : open circles on 3 , shaded to the left. open circle on 5, shaded to the right.